OpenStudy (mathhelp346):

PLEASEE HELP PLEASE!!!: 1. (-1+i √5)^2 2. (2-3i √2)^2

OpenStudy (anonymous):

are they supposed to be equal?

OpenStudy (mathhelp346):

No, they're two seperate problems

OpenStudy (mathhelp346):

Sorry for the confusion

OpenStudy (anonymous):

no problem hang on a sec kk?

OpenStudy (mathhelp346):

Ok, thanks

OpenStudy (anonymous):

the first one should be i*4.84 the second should be i*1.96

OpenStudy (mathhelp346):

How did you figure those out?

OpenStudy (mathhelp346):

I checked the answers at the back of my textbook and it says -4-2i√5. Is that the same thing?

OpenStudy (anonymous):

yes but more complicated(:

OpenStudy (mathhelp346):

Oh, ok. But how did you get those answers anyways?

OpenStudy (anonymous):

step by step lol your book should tell you

OpenStudy (mathhelp346):

It doesn't have an example on this type. I don't know how to solve it with a square root.

OpenStudy (anonymous):

What?

OpenStudy (mathhelp346):

Can you show me how you solved it step by step?

OpenStudy (anikay):

I can!

OpenStudy (mathhelp346):

Ok :o

OpenStudy (anikay):

only cuz of your adorable pikachus though :D

OpenStudy (mathhelp346):

xD Thank you

OpenStudy (anikay):

do you know what \[i^2 \] is?

OpenStudy (mathhelp346):

yes -1

OpenStudy (anikay):

ok, so..... \[(-1+i+\sqrt{5})^2\] that's it?

OpenStudy (mathhelp346):

That's it?

OpenStudy (anikay):

is that the problem in your book?

OpenStudy (mathhelp346):

\[(-1+i \sqrt{5})^{2}\]

OpenStudy (mathhelp346):

It's that

OpenStudy (anikay):

ok thanks

OpenStudy (anikay):

so we square everything inside the parentheses. to get -4

OpenStudy (anikay):

because -1^2 = 1 and i^2 = -1 and sqrt(5)^2 = 5 right?

OpenStudy (mathhelp346):

Yeah

OpenStudy (anikay):

so the answer should be -4.....imo

OpenStudy (anikay):

is there something about the problem I'm missing?

OpenStudy (mathhelp346):

But the answers in my textbook say it's -4-2i√5. Or is your answer and the textbook's the same thing?

OpenStudy (anikay):

they aren't

OpenStudy (anikay):

mine is simplifying the expression..... is that what your book is asking for

OpenStudy (mathhelp346):

The textbook says multiply

OpenStudy (anikay):

....omg I'm so sorry

OpenStudy (anikay):

*facepalm* I'm tired. gonna head to bed after I finish helping you

OpenStudy (anikay):

\[(-1+i \sqrt{5})(-1+i \sqrt{5})\]

OpenStudy (anikay):

foil

OpenStudy (mathhelp346):

1+i√5^2?

OpenStudy (anikay):

ok, F.O.I.L. Front, Outside, Inside, Last

OpenStudy (anikay):

so first we multiply -1 by -1

OpenStudy (anikay):

ok we get 1 next we mutiply \[-1*i \sqrt{5}\]

OpenStudy (anikay):

then we do the same thing again for the inside terms

OpenStudy (anikay):

then the last, which is \[i \sqrt{5}*i \sqrt{5}\]

OpenStudy (anikay):

so all together, we get\[1- i \sqrt{5} - i \sqrt{5} - 5\] which simplifies to \[-4 -2i \sqrt{5}\]

OpenStudy (mathhelp346):

oh ok thank you

OpenStudy (anikay):

so F.O.I.L, gotta remember that :P

OpenStudy (mathhelp346):

Ok ^-^ Good night

OpenStudy (anikay):

here's a picture, just for good measure